OpenStudy (anonymous):

can someone help me find the gcf for 36, 60, and 96

myininaya (myininaya):

36=6*6 60=6*5 96=6*16 what do they all have in common?

OpenStudy (anonymous):

6

myininaya (myininaya):

right!

OpenStudy (anonymous):

yea thanx

myininaya (myininaya):

oops

myininaya (myininaya):

why would you agree with me that 6*5=60?

myininaya (myininaya):

so 6 is the gcf of 36,30,96

myininaya (myininaya):

lol 36=2*3*6 60=2*5*6 96=2*6*8

myininaya (myininaya):

what do they have in common

OpenStudy (anonymous):

is it still 6 cause 6 is in all the problems

myininaya (myininaya):

and what else

OpenStudy (anonymous):

2

myininaya (myininaya):

2*6 is ur gcf=12 is ur gcf

OpenStudy (anonymous):

so im soppose to tims

myininaya (myininaya):

i could have wrote 36=3*12 60=5*12 96=8*12 what do they have in common?

OpenStudy (anonymous):

12

myininaya (myininaya):

yes!

OpenStudy (anonymous):

thanx jus made my momma sit back

myininaya (myininaya):

np